OpenStudy (anonymous):

Find dy/dx: cos(x^2y^2)/sin(x^2y^2)=2

OpenStudy (anonymous):

\[\frac{ dy }{ dx }: \frac{ \cos(x^2y^2) }{ \sin(x^2y^2) }=2x\]

OpenStudy (anonymous):

I switched cos/sin to cot but now i'm lost

myininaya (myininaya):

\[\cot(x^2y^2)=2x\] First let's look at the inside on the left hand side.

myininaya (myininaya):

Do you know how to differentiate x^2y^2 w.r.t x?

OpenStudy (anonymous):

yes I can differentiate x^2y^2

myininaya (myininaya):

ok what do you get when you differentiate that?

myininaya (myininaya):

and then what is the derivative of cot(x) w.r.t x?

myininaya (myininaya):

here i will help you out a little the derivative of cot(x) w.r.t x is -csc^2(x)

myininaya (myininaya):

\[(x^2y^2)'(-\csc^2(x^2y^2))=(2x)'\]

myininaya (myininaya):

let me know if you need more help